OpenStudy (anonymous):

in a survey of 100 people, 30 perferred onions on their hot dogs. what percent pefreed onions?

OpenStudy (angela210793):

30/100 *100=30%

OpenStudy (anonymous):

would it be 3/10?

OpenStudy (anonymous):

30/100 = .3 .3 * 100 = 30% If you're dividing by 100, just use the numerator as your percent After all, that's what a percentage is.

OpenStudy (anonymous):

Yes same thing but 3/10 is a fraction, not a percent 3/10 = .3 .3 * 100 = 30%